# Claim Check-Worthiness Detection: How Well do LLMs Grasp Annotation Guidelines?

### v1

@@ Appendix E Complete prompts
Baseline prompts.
Two naive zero-shot prompts serve as a baseline. For the claim detection task, the baseline is:
Does the following <sentence> contain a factual claim? Answer only with Yes or No.
For the claim check-worthiness detection task, the baseline is:
Does the following <sentence> contain a check-worthy claim? Answer only with Yes or No.
@@
Level Prompt
V1 Categorize the <sentence> spoken in the presidential debates into one of three categories: Non-Factual Sentence (NFS), Unimportant Factual Sentence (UFS) or Check-worthy Factual Sentence (CFS). Use only one of the three labels (NFS, UFS or CFS), do not provide any additional explanation.
V2 Categorize the <sentence> spoken in the presidential debates into three categories: Non-Factual Sentence (NFS): Subjective sentences (opinions, beliefs, declarations) and many questions fall under this category. These sentences do not contain any factual claim. Unimportant Factual Sentence (UFS): These are factual claims but not check-worthy. The general public will not be interested in knowing whether these sentences are true or false. Fact-checkers do not find these sentences as important for checking. Check-worthy Factual Sentence (CFS): They contain factual claims and the general public will be interested in knowing whether the claims are true. Journalists look for these type of claims for fact-checking. Use only one of the three labels (NFS, UFS and CFS), do not provide any additional explanation.
V3 Categorize the <sentence> spoken in the presidential debates into three categories: Non-Factual Sentence (NFS): Subjective sentences (opinions, beliefs, declarations) and many questions fall under this category. These sentences do not contain any factual claim. Here are two such examples. ”But I think it’s time to talk about the future.“ “You remember the last time you said that?” Unimportant Factual Sentence (UFS): These are factual claims but not check-worthy. The general public will not be interested in knowing whether these sentences are true or false. Fact-checkers do not find these sentences as important for checking. Some examples are as follows. “Next Tuesday is Election day.” “Two days ago we ate lunch at a restaurant.” Check-worthy Factual Sentence (CFS): They contain factual claims and the general public will be interested in knowing whether the claims are true. Journalists look for these type of claims for fact-checking. Some examples are: “He voted against the first Gulf War.” “Over a million and a quarter Americans are HIV-positive.” Use only one of the three labels (NFS, UFS and CFS), do not provide any additional explanation.
@@
Level Prompt
V1 Your task is to label the <sentence>. The information I need is whether it is an environmental claim. A broad definition for such a claim is given by the European Commission: Environmental claims refer to the practice of suggesting or otherwise creating the impression that a product or a service is environmentally friendly (i.e., it has a positive impact on the environment) or is less damaging to the environment than competing goods or services. Answer only with Yes or No.
V2 Your task is to label the <sentence>. The information I need is whether it is an environmental claim. A broad definition for such a claim is given by the European Commission: Environmental claims refer to the practice of suggesting or otherwise creating the impression that a product or a service is environmentally friendly (i.e., it has a positive impact on the environment) or is less damaging to the environment than competing goods or services. General principles: You will be presented with a <sentence> and have to decide whether the <sentence> contains an explicit environmental claim. Do not rely on implicit assumptions when you decide on the label. Base your decision on the information that is available within the sentence. However, if a sentence contains an abbreviation, you could consider the meaning of the abbreviation before assigning the label. In case a sentence is too technical/complicated and thus not easily understandable, it usually does not suggest to the average consumer that a product or a service is environmentally friendly and thus can be rejected. Likewise, if a sentence is not specific about having an environmental impact for a product or service, it can be rejected. Answer only with Yes or No.
@@
V3 Your task is to label the <sentence>. The information I need is whether it is an environmental claim. A broad definition for such a claim is given by the European Commission: Environmental claims refer to the practice of suggesting or otherwise creating the impression that a product or a service is environmentally friendly (i.e., it has a positive impact on the environment) or is less damaging to the environment than competing goods or services. General principles: You will be presented with a sentence and have to decide whether the sentence contains an explicit environmental claim. Do not rely on implicit assumptions when you decide on the label. Base your decision on the information that is available within the sentence. However, if a sentence contains an abbreviation, you could consider the meaning of the abbreviation before assigning the label. In case a sentence is too technical/complicated and thus not easily understandable, it usually does not suggest to the average consumer that a product or a service is environmentally friendly and thus can be rejected. Likewise, if a sentence is not specific about having an environmental impact for a product or service, it can be rejected. Examples: <sentence>: Farmers who operate under this scheme are required to dedicate 10% of their land to wildlife preservation. Label: Yes Explanation: Environmental scheme with details on implementation. <sentence>: UPM Biofuels is developing a new feedstock concept by growing Brassica Carinata as a sequential crop in South America. Label: No Explanation: Sentence context would be required to understand whether it is a claim. Answer only with Yes or No, don’t provide any additional explanation.
@@
Level Prompt
V1 A verifiable factual claim is a sentence claiming that something is true, and this can be verified using factual, verifiable information such as statistics, specific examples, or personal testimony. Does the <tweet> contain a verifiable factual claim? Answer only with Yes or No, don’t provide any additional explanation.
V2 A verifiable factual claim is a sentence claiming that something is true, and this can be verified using factual, verifiable information such as statistics, specific examples, or personal testimony. Factual claims include the following: Stating a definition; Mentioning quantity in the present or the past; Making a verifiable prediction about the future; Reference to laws, procedures, and rules of operation; References to images or videos (e.g., "This is a video showing a hospital in Spain.’’); Statements about correlations or causations. Such correlation and causation needs to be explicit, i.e., sentences like "This is why the beaches haven’t closed in Florida.’’ is not a claim because it does not say why explicitly, thus it is not verifiable. Tweets containing personal opinions and preferences are not factual claims. Note: if a tweet is composed of multiple sentences or clauses, at least one full sentence or clause needs to be a claim in order for the tweet to contain a factual claim. If a claim exist in a sub-sentence or sub-clause then tweet is not considered to have a factual claim. For example, "My new favorite thing is Italian mayors and regional presidents LOSING IT at people violating quarantine’’ is not a claim, however, it is an opinion. Moreover, if we consider "Italian mayors and regional presidents LOSING IT at people violating quarantine’’ it would be a claim. In addition, when answering this question, annotator should not open the tweet URL. Does the <tweet> contain a verifiable factual claim? Answer only with Yes or No.
@@
V3 A verifiable factual claim is a sentence claiming that something is true, and this can be verified using factual, verifiable information such as statistics, specific examples, or personal testimony. Factual claims include the following: Stating a definition; Mentioning quantity in the present or the past; Making a verifiable prediction about the future; Reference to laws, procedures, and rules of operation; References to images or videos (e.g., "This is a video showing a hospital in Spain.’’); Statements about correlations or causations. Such correlation and causation needs to be explicit, i.e., sentences like "This is why the beaches haven’t closed in Florida.’’ is not a claim because it does not say why explicitly, thus it is not verifiable. Tweets containing personal opinions and preferences are not factual claims. Note: if a tweet is composed of multiple sentences or clauses, at least one full sentence or clause needs to be a claim in order for the tweet to contain a factual claim. If a claim exist in a sub-sentence or sub-clause then tweet is not considered to have a factual claim. For example, "My new favorite thing is Italian mayors and regional presidents LOSING IT at people violating quarantine’’ is not a claim, however, it is an opinion. Moreover, if we consider "Italian mayors and regional presidents LOSING IT at people violating quarantine’’ it would be a claim. In addition, when answering this question, annotator should not open the tweet URL. Does the <tweet> contain a verifiable factual claim? Answer only with Yes or No. Examples: Tweet: Please don’t take hydroxychloroquine (Plaquenil) plus Azithromycin for #COVID19 UNLESS your doctor prescribes it. Both drugs affect the QT interval of your heart and can lead to arrhythmias and sudden death, especially if you are taking other meds or have a heart condition. Label: Yes Explanation: There is a claim in the text. Tweet: Saw this on Facebook today and it’s a must read for all those idiots clearing the shelves #coronavirus #toiletpapercrisis #auspol Label: No Explanation: There is no claim in the text. Answer only with Yes or No, don’t provide any additional explanation.
@@
Level Prompt
V1 It is important that a verifiable factual check-worthy claim be verified by a professional fact-checker, as the claim may cause harm to society, specific person(s), company(s), product(s), or some government entities. However, not all factual claims are important or worth fact-checking by a professional fact-checker, as this very time-consuming. Do you think that a professional fact-checker should verify the claim in the <tweet>? Labels: No, no need to check; No, too trivial to check; Yes, not urgent; Yes, very urgent. Decide on one label. Then, answer only with Yes or No.
V2 It is important that a verifiable factual check-worthy claim be verified by a professional fact-checker, as the claim may cause harm to society, specific person(s), company(s), product(s), or some government entities. However, not all factual claims are important or worth fact-checking by a professional fact-checker, as this very time-consuming. Do you think that a professional fact-checker should verify the claim in the <tweet>? Labels: No, no need to check: the tweet does not need to be fact-checked, e.g., be- cause it is not interesting, a joke, or does not contain any claim. No, too trivial to check: the tweet is worth fact-checking, however, this does not require a professional fact-checker, i.e., a non-expert might be able to fact-check the claim. For example, one can verify the information using reliable sources such as the official website of the WHO, etc. An example of a claim is as follows: “The GDP of the USA grew by 50% last year.” Yes, not urgent: the tweet should be fact-checked by a professional fact-checker, however, this is not urgent or critical; Yes, very urgent: the tweet can cause immediate harm to a large number of people; therefore, it should be verified as soon as possible by a professional fact-checker; Decide on one label. Then, answer only with Yes or No.
@@
V3 It is important to verify a factual claim by a professional fact-checker, which can cause harm to the society, specific person(s), company(s), product(s) or government entities. However, not all factual claims are important or worthwhile to be fact-checked by a professional fact-checker as it is a time-consuming procedure. Do you think that a professional fact-checker should verify the claim in the <tweet>? Labels: No, no need to check: the tweet does not need to be fact-checked, e.g., be- cause it is not interesting, a joke, or does not contain any claim. No, too trivial to check: the tweet is worth fact-checking, however, this does not require a professional fact-checker, i.e., a non-expert might be able to fact-check the claim. For example, one can verify the information using reliable sources such as the official website of the WHO, etc. An example of a claim is as follows: “The GDP of the USA grew by 50% last year.” Yes, not urgent: the tweet should be fact-checked by a professional fact-checker, however, this is not urgent or critical; Yes, very urgent: the tweet can cause immediate harm to a large number of people; therefore, it should be verified as soon as possible by a professional fact-checker; Examples: Tweet: ’Wash your hands like you’ve been chopping jalapeños and need to change a contact lens’ says BC Public Health Officer Dr. Bonnie Henry re. ways to protect against #coronavirus #Covid_19 Label: Yes, not urgent Explanation: Overall it is less important for a professional fact-checker to verify this information. The statement does not harm anyone. The truth value of whether the official said the statement is not important. Also it appears that washing hands is very important to protect oneself from the virus. Tweet: ALERT! The corona virus can be spread through internationaly printed albums. If you have any albums at home, put on some gloves, put all the albums in a box and put it outside the front door tonight. I’m collecting all the boxes tonight for safety. Think of your health. Label: No, no need to check Explanation: This is joke and no need to check by a professional fact checker. Decide on one label. Then, answer only with Yes or No.
@@
Level Prompt
V1 The task is to select verifiable statements from political speeches for fact-checking. Given a <statement> from a political speech, answer the question. Does the <statement> explicitly present any verifiable factual information? Answer with A, B or C only. A - Yes, B - Maybe, C - No.
V2 The task is to select verifiable statements from political speeches for fact-checking. Given a <statement> from a political speech, answer the question following the guidelines. Definitions and guidelines: Fact: A fact is a statement or assertion that can be objectively verified as true or false based on empirical evidence or reality. Opinion: An opinion is a judgment based on facts, an attempt to draw a reasonable conclusion from factual evidence. While the underlying facts can be verified, the derived opinion remains subjective and is not universally verifiable. Context: Make sure to consider a small context of the target statement (the previous and next sentence) when annotating. Some statements require context to understand the meaning. Factual claim: A factual claim is a statement that explicitly presents some verifiable facts. Statements with subjective components like opinions can also be factual claims if they explicitly present objectively verifiable facts. Opinion with Facts: Opinions can also be based on factual information. When does an opinion explicitly present a fact: Many opinions are more or less based on some factual information. However, some facts are explicitly presented by the speakers, while others are not. What is verifiable: The verifiability of the factual information depends on how specific it is. If there is enough specific information to guide a general fact-checker in checking it, the factual information is verifiable. Otherwise, it is not verifiable. The question: Does the <statement> explicitly present any verifiable factual information? Answer with A, B or C only. A - Yes, the statement contains factual information with enough specific details that a fact-checker knows how to verify it. E.g., Birmingham is small in population compared to London. B - Maybe, the statement seems to contain some factual information. However, there are certain ambiguities (e.g., lack of specificity) making it hard to determine the verifiability. E.g., Birmingham is small compared to London. (lack of details about what standard Birmingham is small) C - No, the statement contains no verifiable factual information. Even if there is some, it is clearly unverifiable. E.g., Birmingham is small.
@@
V3 The task is to select verifiable statements from political speeches for fact-checking. Given a statement from a political speech and its context, answer the question following the guidelines. Definitions and guidelines: Fact: A fact is a statement or assertion that can be objectively verified as true or false based on empirical evidence or reality. Opinion: An opinion is a judgment based on facts, an attempt to draw a reasonable conclusion from factual evidence. While the underlying facts can be verified, the derived opinion remains subjective and is not universally verifiable. Factual claim: A factual claim is a statement that explicitly presents some verifiable facts. Statements with subjective components like opinions can also be factual claims if they explicitly present objectively verifiable facts. Context: Make sure to consider a small context of the target statement (the previous and next sentence) when annotating. Some statements require context to understand the meaning. For example: E1. “... Just consider what we did last year for the middle class in California, sending 12 billion dollars back – the largest state tax rebate in American history. <statement> But we didn’t stop there. <_statement> We raised the minimum wage. We increased paid sick leave. Provided more paid family leave. Expanded child care to help working parents...” Without the context, the sentence marked with <statement> seems an incomplete sentence. With the context, we know the speaker is claiming a bunch of verifiable achievements of their administration. E2. “... When I first stood before this chamber three years ago, I declared war on criminals and asked for the Legislature to repeal and replace the catch-and-release policies in SB 91. <statement> With the help of many of you, we got it done. <_statement> Policies do matter. We’ve seen our overall crime rate decline by 10 percent in 2019 and another 18.5 percent in 2020! ...” The part marked with <statement> claims that the policies against crimes have been “done”, which is verifiable. It needs context to understand it. Opinion with Facts: Opinions can also be based on factual information. For example: E1. “I am proud to report that on top of the local improvements, the state has administered projects in almost all 67 counties already, and like I said, we’ve only just begun.” The speaker’s “proud of” is a subjective opinion. However, the content of pride (administered projects) is factual information. E2. “I first want to thank my wife of 34 years, First Lady Rose Dunleavy.” The speaker expresses their thankfulness to their wife. However, there is factual information about the first lady’s name and the length of their marriage. When does an opinion explicitly present a fact: Many opinions are more or less based on some factual information. However, some facts are explicitly presented by the speakers, while others are not. Explicit presentation means the fact is directly entailed by the opinion without extrapolation: E1. “The pizza is delicious.” This opinion seems to be based on the fact that “pizza is a kind of food”. However, this fact is not explicitly presented. E2. “I first want to thank my wife of 34 years, First Lady Rose Dunleavy.” The name of the speaker’s wife and their year of marriage are explicitly presented. What is verifiable: The verifiability of the factual information depends on how specific it is. If there is enough specific information to guide a general fact-checker in checking it, the factual information is verifiable. Otherwise, it is not verifiable. E1. “Birmingham is small.” is not verifiable because it lacks any specific information for determining veracity. It leans more toward subjective opinion. E2. “Birmingham is small, compared to London” is more verifiable than E1. A fact-checker can retrieve the city size, population size ...etc., of London and Birmingham to compare them. However, what to compare to prove Birmingham’s “small” is not specific enough. E3. “Birmingham is small in population size, compared to London” is more verifiable than E1 and E2. A fact-checker now knows it is exactly the population size to be compared. The question: Does the <statement> explicitly present any verifiable factual information? Answer with A, B or C only. A - Yes, the statement contains factual information with enough specific details that a fact-checker knows how to verify it. E.g., Birmingham is small in population compared to London. B - Maybe, Maybe, the statement seems to contain some factual information. However, there are certain ambiguities (e.g., lack of specificity) making it hard to determine the verifiability. E.g., Birmingham is small compared to London. (lack of details about what standard Birmingham is small) C - No, the statement contains no verifiable factual information. Even if there is some, it is clearly unverifiable. E.g., Birmingham is small.

### v2

@@ Appendix D Complete prompts
Level Prompt V1 Categorize the <sentence> spoken in the presidential debates into one of three categories: Non-Factual Sentence (NFS), Unimportant Factual Sentence (UFS) or Check-worthy Factual Sentence (CFS). Use only one of the three labels (NFS, UFS or CFS), do not provide any additional explanation. V2 Categorize the <sentence> spoken in the presidential debates into three categories: Non-Factual Sentence (NFS): Subjective sentences (opinions, beliefs, declarations) and many questions fall under this category. These sentences do not contain any factual claim. Unimportant Factual Sentence (UFS): These are factual claims but not check-worthy. The general public will not be interested in knowing whether these sentences are true or false. Fact-checkers do not find these sentences as important for checking. Check-worthy Factual Sentence (CFS): They contain factual claims and the general public will be interested in knowing whether the claims are true. Journalists look for these type of claims for fact-checking. Use only one of the three labels (NFS, UFS and CFS), do not provide any additional explanation. V3 Categorize the <sentence> spoken in the presidential debates into three categories: Non-Factual Sentence (NFS): Subjective sentences (opinions, beliefs, declarations) and many questions fall under this category. These sentences do not contain any factual claim. Here are two such examples. ”But I think it’s time to talk about the future.“ “You remember the last time you said that?” Unimportant Factual Sentence (UFS): These are factual claims but not check-worthy. The general public will not be interested in knowing whether these sentences are true or false. Fact-checkers do not find these sentences as important for checking. Some examples are as follows. “Next Tuesday is Election day.” “Two days ago we ate lunch at a restaurant.” Check-worthy Factual Sentence (CFS): They contain factual claims and the general public will be interested in knowing whether the claims are true. Journalists look for these type of claims for fact-checking. Some examples are: “He voted against the first Gulf War.” “Over a million and a quarter Americans are HIV-positive.” Use only one of the three labels (NFS, UFS and CFS), do not provide any additional explanation.
@@
Level Prompt V1 Your task is to label the <sentence>. The information I need is whether it is an environmental claim. A broad definition for such a claim is given by the European Commission: Environmental claims refer to the practice of suggesting or otherwise creating the impression that a product or a service is environmentally friendly (i.e., it has a positive impact on the environment) or is less damaging to the environment than competing goods or services. Answer only with Yes or No. V2 Your task is to label the <sentence>. The information I need is whether it is an environmental claim. A broad definition for such a claim is given by the European Commission: Environmental claims refer to the practice of suggesting or otherwise creating the impression that a product or a service is environmentally friendly (i.e., it has a positive impact on the environment) or is less damaging to the environment than competing goods or services. General principles: You will be presented with a <sentence> and have to decide whether the <sentence> contains an explicit environmental claim. Do not rely on implicit assumptions when you decide on the label. Base your decision on the information that is available within the sentence. However, if a sentence contains an abbreviation, you could consider the meaning of the abbreviation before assigning the label. In case a sentence is too technical/complicated and thus not easily understandable, it usually does not suggest to the average consumer that a product or a service is environmentally friendly and thus can be rejected. Likewise, if a sentence is not specific about having an environmental impact for a product or service, it can be rejected. Answer only with Yes or No. V3 Your task is to label the <sentence>. The information I need is whether it is an environmental claim. A broad definition for such a claim is given by the European Commission: Environmental claims refer to the practice of suggesting or otherwise creating the impression that a product or a service is environmentally friendly (i.e., it has a positive impact on the environment) or is less damaging to the environment than competing goods or services. General principles: You will be presented with a sentence and have to decide whether the sentence contains an explicit environmental claim. Do not rely on implicit assumptions when you decide on the label. Base your decision on the information that is available within the sentence. However, if a sentence contains an abbreviation, you could consider the meaning of the abbreviation before assigning the label. In case a sentence is too technical/complicated and thus not easily understandable, it usually does not suggest to the average consumer that a product or a service is environmentally friendly and thus can be rejected. Likewise, if a sentence is not specific about having an environmental impact for a product or service, it can be rejected. Examples: <sentence>: Farmers who operate under this scheme are required to dedicate 10% of their land to wildlife preservation. Label: Yes Explanation: Environmental scheme with details on implementation. <sentence>: UPM Biofuels is developing a new feedstock concept by growing Brassica Carinata as a sequential crop in South America. Label: No Explanation: Sentence context would be required to understand whether it is a claim. Answer only with Yes or No, don’t provide any additional explanation.
@@
Level Prompt V1 A verifiable factual claim is a sentence claiming that something is true, and this can be verified using factual, verifiable information such as statistics, specific examples, or personal testimony. Does the <tweet> contain a verifiable factual claim? Answer only with Yes or No, don’t provide any additional explanation. V2 A verifiable factual claim is a sentence claiming that something is true, and this can be verified using factual, verifiable information such as statistics, specific examples, or personal testimony. Factual claims include the following: Stating a definition; Mentioning quantity in the present or the past; Making a verifiable prediction about the future; Reference to laws, procedures, and rules of operation; References to images or videos (e.g., "This is a video showing a hospital in Spain.’’); Statements about correlations or causations. Such correlation and causation needs to be explicit, i.e., sentences like "This is why the beaches haven’t closed in Florida.’’ is not a claim because it does not say why explicitly, thus it is not verifiable. Tweets containing personal opinions and preferences are not factual claims. Note: if a tweet is composed of multiple sentences or clauses, at least one full sentence or clause needs to be a claim in order for the tweet to contain a factual claim. If a claim exist in a sub-sentence or sub-clause then tweet is not considered to have a factual claim. For example, "My new favorite thing is Italian mayors and regional presidents LOSING IT at people violating quarantine’’ is not a claim, however, it is an opinion. Moreover, if we consider "Italian mayors and regional presidents LOSING IT at people violating quarantine’’ it would be a claim. In addition, when answering this question, annotator should not open the tweet URL. Does the <tweet> contain a verifiable factual claim? Answer only with Yes or No. V3 A verifiable factual claim is a sentence claiming that something is true, and this can be verified using factual, verifiable information such as statistics, specific examples, or personal testimony. Factual claims include the following: Stating a definition; Mentioning quantity in the present or the past; Making a verifiable prediction about the future; Reference to laws, procedures, and rules of operation; References to images or videos (e.g., "This is a video showing a hospital in Spain.’’); Statements about correlations or causations. Such correlation and causation needs to be explicit, i.e., sentences like "This is why the beaches haven’t closed in Florida.’’ is not a claim because it does not say why explicitly, thus it is not verifiable. Tweets containing personal opinions and preferences are not factual claims. Note: if a tweet is composed of multiple sentences or clauses, at least one full sentence or clause needs to be a claim in order for the tweet to contain a factual claim. If a claim exist in a sub-sentence or sub-clause then tweet is not considered to have a factual claim. For example, "My new favorite thing is Italian mayors and regional presidents LOSING IT at people violating quarantine’’ is not a claim, however, it is an opinion. Moreover, if we consider "Italian mayors and regional presidents LOSING IT at people violating quarantine’’ it would be a claim. In addition, when answering this question, annotator should not open the tweet URL. Does the <tweet> contain a verifiable factual claim? Answer only with Yes or No. Examples: Tweet: Please don’t take hydroxychloroquine (Plaquenil) plus Azithromycin for #COVID19 UNLESS your doctor prescribes it. Both drugs affect the QT interval of your heart and can lead to arrhythmias and sudden death, especially if you are taking other meds or have a heart condition. Label: Yes Explanation: There is a claim in the text. Tweet: Saw this on Facebook today and it’s a must read for all those idiots clearing the shelves #coronavirus #toiletpapercrisis #auspol Label: No Explanation: There is no claim in the text. Answer only with Yes or No, don’t provide any additional explanation.
@@
Level Prompt V1 It is important that a verifiable factual check-worthy claim be verified by a professional fact-checker, as the claim may cause harm to society, specific person(s), company(s), product(s), or some government entities. However, not all factual claims are important or worth fact-checking by a professional fact-checker, as this very time-consuming. Do you think that a professional fact-checker should verify the claim in the <tweet>? Labels: No, no need to check; No, too trivial to check; Yes, not urgent; Yes, very urgent. Decide on one label. Then, answer only with Yes or No. V2 It is important that a verifiable factual check-worthy claim be verified by a professional fact-checker, as the claim may cause harm to society, specific person(s), company(s), product(s), or some government entities. However, not all factual claims are important or worth fact-checking by a professional fact-checker, as this very time-consuming. Do you think that a professional fact-checker should verify the claim in the <tweet>? Labels: No, no need to check: the tweet does not need to be fact-checked, e.g., be- cause it is not interesting, a joke, or does not contain any claim. No, too trivial to check: the tweet is worth fact-checking, however, this does not require a professional fact-checker, i.e., a non-expert might be able to fact-check the claim. For example, one can verify the information using reliable sources such as the official website of the WHO, etc. An example of a claim is as follows: “The GDP of the USA grew by 50% last year.” Yes, not urgent: the tweet should be fact-checked by a professional fact-checker, however, this is not urgent or critical; Yes, very urgent: the tweet can cause immediate harm to a large number of people; therefore, it should be verified as soon as possible by a professional fact-checker; Decide on one label. Then, answer only with Yes or No. V3 It is important to verify a factual claim by a professional fact-checker, which can cause harm to the society, specific person(s), company(s), product(s) or government entities. However, not all factual claims are important or worthwhile to be fact-checked by a professional fact-checker as it is a time-consuming procedure. Do you think that a professional fact-checker should verify the claim in the <tweet>? Labels: No, no need to check: the tweet does not need to be fact-checked, e.g., be- cause it is not interesting, a joke, or does not contain any claim. No, too trivial to check: the tweet is worth fact-checking, however, this does not require a professional fact-checker, i.e., a non-expert might be able to fact-check the claim. For example, one can verify the information using reliable sources such as the official website of the WHO, etc. An example of a claim is as follows: “The GDP of the USA grew by 50Yes, not urgent: the tweet should be fact-checked by a professional fact-checker, however, this is not urgent or critical; Yes, very urgent: the tweet can cause immediate harm to a large number of people; therefore, it should be verified as soon as possible by a professional fact-checker; Examples: Tweet: Wash your hands like you’ve been chopping jalapeños and need to change a contact lens’’ says BC Public Health Officer Dr. Bonnie Henry re. ways to protect against #coronavirus #Covid_19 Label: Yes, not urgent Explanation: Overall it is less important for a professional fact-checker to verify this information. The statement does not harm anyone. The truth value of whether the official said the statement is not important. Also it appears that washing hands is very important to protect oneself from the virus. Tweet: ALERT! The corona virus can be spread through internationaly printed albums. If you have any albums at home, put on some gloves, put all the albums in a box and put it outside the front door tonight. I’m collecting all the boxes tonight for safety. Think of your health. Label: No, no need to check Explanation: This is joke and no need to check by a professional fact checker. Decide on one label. Then, answer only with Yes or No.
@@
Level Prompt V1 The task is to select verifiable statements from political speeches for fact-checking. Given a <statement >from a political speech, answer the question. Does the <statement >explicitly present any verifiable factual information? Answer with A, B or C only. A - Yes, B - Maybe, C - No. V2 The task is to select verifiable statements from political speeches for fact-checking. Given a <statement >from a political speech, answer the question following the guidelines. Definitions and guidelines: Fact: A fact is a statement or assertion that can be objectively verified as true or false based on empirical evidence or reality. Opinion: An opinion is a judgment based on facts, an attempt to draw a reasonable conclusion from factual evidence. While the underlying facts can be verified, the derived opinion remains subjective and is not universally verifiable. Context: Make sure to consider a small context of the target statement (the previous and next sentence) when annotating. Some statements require context to understand the meaning. Factual claim: A factual claim is a statement that explicitly presents some verifiable facts. Statements with subjective components like opinions can also be factual claims if they explicitly present objectively verifiable facts. Opinion with Facts: Opinions can also be based on factual information. When does an opinion explicitly present a fact: Many opinions are more or less based on some factual information. However, some facts are explicitly presented by the speakers, while others are not. What is verifiable: The verifiability of the factual information depends on how specific it is. If there is enough specific information to guide a general fact-checker in checking it, the factual information is verifiable. Otherwise, it is not verifiable. The question: Does the <statement >explicitly present any verifiable factual information? Answer with A, B or C only. A - Yes, the statement contains factual information with enough specific details that a fact-checker knows how to verify it. E.g., Birmingham is small in population compared to London. B - Maybe, the statement seems to contain some factual information. However, there are certain ambiguities (e.g., lack of specificity) making it hard to determine the verifiability. E.g., Birmingham is small compared to London. (lack of details about what standard Birmingham is small) C - No, the statement contains no verifiable factual information. Even if there is some, it is clearly unverifiable. E.g., Birmingham is small.
@@
Level Prompt V3 The task is to select verifiable statements from political speeches for fact-checking. Given a statement from a political speech and its context, answer the question following the guidelines. Definitions and guidelines: Fact: A fact is a statement or assertion that can be objectively verified as true or false based on empirical evidence or reality. Opinion: An opinion is a judgment based on facts, an attempt to draw a reasonable conclusion from factual evidence. While the underlying facts can be verified, the derived opinion remains subjective and is not universally verifiable. Factual claim: A factual claim is a statement that explicitly presents some verifiable facts. Statements with subjective components like opinions can also be factual claims if they explicitly present objectively verifiable facts. Context: Make sure to consider a small context of the target statement (the previous and next sentence) when annotating. Some statements require context to understand the meaning. For example: E1. “... Just consider what we did last year for the middle class in California, sending 12 billion dollars back – the largest state tax rebate in American history. <statement >But we didn’t stop there. <_statement >We raised the minimum wage. We increased paid sick leave. Provided more paid family leave. Expanded child care to help working parents...” Without the context, the sentence marked with <statement >seems an incomplete sentence. With the context, we know the speaker is claiming a bunch of verifiable achievements of their administration. E2. “... When I first stood before this chamber three years ago, I declared war on criminals and asked for the Legislature to repeal and replace the catch-and-release policies in SB 91. <statement >With the help of many of you, we got it done. <_statement> Policies do matter. We’ve seen our overall crime rate decline by 10 percent in 2019 and another 18.5 percent in 2020! ...” The part marked with <statement> claims that the policies against crimes have been “done”, which is verifiable. It needs context to understand it. Opinion with Facts: Opinions can also be based on factual information. For example: E1. “I am proud to report that on top of the local improvements, the state has administered projects in almost all 67 counties already, and like I said, we’ve only just begun.” The speaker’s “proud of” is a subjective opinion. However, the content of pride (administered projects) is factual information. E2. “I first want to thank my wife of 34 years, First Lady Rose Dunleavy.” The speaker expresses their thankfulness to their wife. However, there is factual information about the first lady’s name and the length of their marriage. When does an opinion explicitly present a fact: Many opinions are more or less based on some factual information. However, some facts are explicitly presented by the speakers, while others are not. Explicit presentation means the fact is directly entailed by the opinion without extrapolation: E1. “The pizza is delicious.” This opinion seems to be based on the fact that “pizza is a kind of food”. However, this fact is not explicitly presented. E2. “I first want to thank my wife of 34 years, First Lady Rose Dunleavy.” The name of the speaker’s wife and their year of marriage are explicitly presented. What is verifiable: The verifiability of the factual information depends on how specific it is. If there is enough specific information to guide a general fact-checker in checking it, the factual information is verifiable. Otherwise, it is not verifiable. E1. “Birmingham is small.” is not verifiable because it lacks any specific information for determining veracity. It leans more toward subjective opinion. E2. “Birmingham is small, compared to London” is more verifiable than E1. A fact-checker can retrieve the city size, population size ...etc., of London and Birmingham to compare them. However, what to compare to prove Birmingham’s “small” is not specific enough. E3. “Birmingham is small in population size, compared to London” is more verifiable than E1 and E2. A fact-checker now knows it is exactly the population size to be compared. The question: Does the <statement> explicitly present any verifiable factual information? Answer with A, B or C only. A - Yes, the statement contains factual information with enough specific details that a fact-checker knows how to verify it. E.g., Birmingham is small in population compared to London. B - Maybe, Maybe, the statement seems to contain some factual information. However, there are certain ambiguities (e.g., lack of specificity) making it hard to determine the verifiability. E.g., Birmingham is small compared to London. (lack of details about what standard Birmingham is small) C - No, the statement contains no verifiable factual information. Even if there is some, it is clearly unverifiable. E.g., Birmingham is small.